1,3,5,7,9,11,13,15,17,19 ai sonkha guli diye 50 milao plz.
Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

1,3,5,7,9,11,13,15,17,19
:
यह एक अंकगणितीय अनुक्रम है, आम अंतर 2 है, इसलिए नतीज शब्द है
:
x(n) = 1 + d(n-1)
:
अगले शब्द x (11) है
:
x(11) = 1 + 2(11 -1) = 21
:

Answer by ikleyn(52908)

.
There is a bunch of lessons on arithmetic progressions in this site:
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
    - Mathematical induction and arithmetic progressions
    - One characteristic property of arithmetic progressions
    - Solved problems on arithmetic progressions


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Arithmetic progressions".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.